34) A zero-coupon bond with a 15-year maturity and a face value of $1,000 is sold at $123. What is the bond's yield to maturity?

A. about 15%

B. about 18%

C. about 25%

D. about 26%

E. about 27%
Answer #1

Current value=Face value/(1+YTM)^time period

123=1000/(1+YTM)^15

(1+YTM)^15=(1000/123)

1+YTM=(1000/123)^(1/15)

YTM=1.15-1

=15%(Approx).